# Finite and full scale localization for the multi-frequency quasi-periodic CMV matrices

### v2

@@ 3.2 Estimation of the Lyapunov Exponent
It is straightforward to verify that to verify that $M(\omega,z;x)$ is conjugate to an $SL(2,\mathbb{R})$ matrix $T(\omega,z;x)$ via Q i.e.,
$Q=-\frac{1}{1+i}\left[\begin{array}[]{cc}1&-i\\ 1&i\end{array}\right]\in\mathbb{U}(2),$
@@
$T(\omega,z;x)=Q^{*}M(\omega,z;x)Q\in SL(2,\mathbb{R}),$
@@
We can finally prove this lemma by summing over $2^{k}n$ ∎
@@
$M_{k,n}=\frac{1}{\rho(x_{k}+n\omega_{k})}\left[\begin{array}[]{cc}\sqrt{z_{k}}&-\frac{\overline{\alpha}(x_{k}+n\omega_{k})}{\sqrt{z_{k}}}\\ -\alpha(x_{k}+n\omega_{k})&1/\sqrt{z_{k}}\end{array}\right],\quad k=1,2.$
@@
It’s not difficult to check that $M_{k,n}$ is conjugate to an $SL(2,\mathbb{R})$ matrix $T_{k,n}$ through $Q$, that is,
@@
According to the polar decomposition of $SL(2,\mathbb{C})$ matrices,
@@
for any $x_{0}\in\mathbb{T}^{d}\backslash\mathcal{B}_{n,\omega_{0},z_{0}}$ and $(\omega,z,x)\in\mathbb{C}^{d}\times\partial\mathbb{D}\times\mathbb{C}^{d}$ such that
@@
provided that $\sigma A\geq 1$. Take $x_{0}\in\mathbb{T}^{d}\backslash\mathcal{B}_{n,\omega_{0},z_{0}}$. Applying (AP) with $\mu=\exp(l\gamma/2)$, we have
$\Big{|}\log\|M_{m}(\omega_{0},z_{0};x_{0})\|+\sum_{j=2}^{m-1}\log\|B_{j}(\omega_{0},z_{0};x_{0})\|-\sum_{j=1}^{m-1}\log\|B_{j+1}(\omega_{0},z_{0};x_{0})B_{j}(\omega_{0},z_{0};x_{0})\|\Big{|}$
@@
$\big{|}\log\|M_{m}(\omega,z;x)\|-\log\|M_{m}(\omega_{0},z_{0};x_{0})\|\big{|}$ $\leq\Big{(}\frac{4}{\sqrt{1-\|\alpha\|}_{h}}\Big{)}^{n}(|\omega-\omega_{0}|+|z-z_{0}|+|x-x_{0}|)$
@@
[13, Lemma 4.1] Let $10<\rho<$ 1 and suppose $f$ is subharmonic on
@@
[…]
@@
[…]
@@ 3.3 LDT Estimate for the Entries of the Monodromy
$\Theta_{n}=\left[\begin{array}[]{cc}\overline{\alpha}_{n}&\rho_{n}\\ \rho_{n}&-\alpha_{n}\end{array}\right].$
@@
$\mathcal{L}_{+}=\left[\begin{matrix}\Theta_{0}&~{}&~{}\\ ~{}&\Theta_{2}&~{}\\ ~{}&~{}&\ddots\end{matrix}\right],\quad\mathcal{M}_{+}=\left[\begin{matrix}\mathbf{1}&~{}&~{}\\ ~{}&\Theta_{1}&~{}\\ ~{}&~{}&\ddots\end{matrix}\right].$
@@
Without loss of generality, we restrict the extended CMV matrix to the interval $[0,n-1]$. We denote the characteristic determinant of matrix [$\mathcal{E}^{\beta,\eta}_{[0,n-1]}$ as $\varphi^{\beta,\eta}_{[0,n-1]}(\omega,z;x)=\det\big{(}z-\mathcal{E}^{\beta,\eta}_{[0,n-1]}\big{)}$. According to the results in [26, Theorem 2], the relation between this characteristic determinant and the $n$-step transfer matrix is
@@
Let $N_{0}$ be as in Lemma 3.15 and $10<\theta<$ 1. Fix some $z$ and let $\varphi_{[0,n-1]}^{\beta,\eta}=\det\big{(}z-\mathcal{E}^{\beta,\eta}_{[0,n-1]}\big{)}$. Then for any $N_{0}\leq n<N^{\frac{1-\theta}{2}}$,
@@
Now we introduce the properties of matrices in $SL(2,\mathbb{C})$. For any $A\in SL(2,\mathbb{C})$, there are unit vectors $\underline{u}^{+}_{A}$, $\underline{u}^{-}_{A}$, $\underline{v}^{+}_{A}$, $\underline{v}^{-}_{A}$ such that $A\underline{u}^{+}_{A}=\|A\|\underline{v}^{+}_{A}$, $A\underline{u}^{-}_{A}=\|A\|^{-1}\underline{v}^{-}_{A}$. Moreover, $\underline{u}^{+}_{A}\bot\underline{u}^{-}_{A}$ and $\underline{v}^{+}_{A}\bot\underline{v}^{-}_{A}$. In what follows, we denote that $a\wedge b:=\min\{a,b\}$ for positive numbers $a,b$ and $\varphi\wedge\psi$ stands for the exterior product of vectors $\varphi,\psi$.
@@
From the Lemma 3.2, we have
@@
From (3.19), the transfer matrix $M_{n}(\omega,z;x)\in SU(1,1)$, then we can verify that
$T_{n}(\omega,z;x)=Q^{*}M_{n}(\omega,z;x)Q=(\sqrt{z})^{-n}\Big{(}\prod_{j=0}^{n-1}\frac{1}{\rho_{j}}\Big{)}\left(\begin{matrix}t_{1}&t_{2}\\ t_{3}&t_{4}\end{matrix}\right)\in SL(2,\mathbb{R}),$ (3.26)
@@
$\mathcal{P}_{n-1}=\left[\begin{array}[]{ccccccc}-\overline{\alpha}_{1}\rho_{0}&-\rho_{1}\rho_{0}&&&&\\ -\overline{\alpha}_{2}\rho_{1}&z+\overline{\alpha}_{2}\alpha_{1}&-\overline{\alpha}_{3}\rho_{2}&-\rho_{3}\rho_{2}&&\\ -\rho_{2}\rho_{1}&\rho_{2}\alpha_{1}&z+\overline{\alpha}_{3}\alpha_{2}&\rho_{3}\alpha_{2}&&\\ &&-\overline{\alpha}_{4}\rho_{3}&z+\overline{\alpha}_{4}\alpha_{3}&&\\ &&&&\ddots&\\ &&&&&z+\overline{\eta}\alpha_{n-2}\end{array}\right].$
@@
[…]
@@
[…]
@@
([$\|(\mathcal{E}^{\beta,\eta}_{[0,n-1]}(\omega,z;x)-z)^{-1}\|\geq C\exp(H+n^{\nu}/d)$
@@
Due to the structure of the matrix [$\mathcal{E}^{\beta,\eta}_{[0,n-1]}$, there exists some constant $C$ such that
[$\|\mathcal{E}^{\beta,\eta}_{[0,n-1]}(\omega,z;x)-\mathcal{E}^{\beta,\eta}_{[0,n-1]}(\omega,z;x^{\prime})\|\leq C|x-x^{\prime}|.$
@@
([$\|(\mathcal{E}^{\beta,\eta}_{[0,n-1]}(\omega,z;x^{\prime})-E)^{-1}\|\|\mathcal{E}^{\beta,\eta}_{[0,n-1]}(\omega,z;x)-\mathcal{E}^{\beta,\eta}_{[0,n-1]}(\omega,z;x^{\prime})\|<1,$
@@
([$\|(\mathcal{E}^{\beta,\eta}_{[0,n-1]}(\omega,z;x^{\prime})-E)^{-1}\|\|\mathcal{E}^{\beta,\eta}_{[0,n-1]}(\omega,z;x)-\mathcal{E}^{\beta,\eta}_{[0,n-1]}(\omega,z;x^{\prime})\|$
=([$=\|(\mathcal{E}^{\beta,\eta}_{[0,n-1]}(\omega,z;x^{\prime})-E)^{-1}\|\|(\mathcal{E}^{\beta,\eta}_{[0,n-1]}(\omega,z;x)-E)-(\mathcal{E}^{\beta,\eta}_{[0,n-1]}(\omega,z;x^{\prime})-E)\|$
@@
Thus, $E-[E-(\mathcal{E}^{\beta,\eta}_{[0,n-1]}(\omega,z;x^{\prime})-E)^{-1}(\mathcal{E}^{\beta,\eta}_{[0,n-1]}(\omega,z;x)-E)]$ is invertible and this implies that [$\mathcal{E}^{\beta,\eta}_{[0,n-1]}(\omega,z;x)-E$ is invertible. Then the second statement follows. ∎
@@ 3.4 Poisson Formula for CMV Matrix
$u(m)=$ $G^{\beta,\eta}_{[a,b]}(a,m;z)\begin{cases}(z\overline{\beta}-\alpha_{a})u(a)-\rho_{a}u(a+1),\quad&a\text{ even,}\\ (z\alpha_{a}-\beta)u(a)+z\rho_{a}u(a+1),&a\text{ odd,}\end{cases}$
$+G^{\beta,\eta}_{[a,b]}(m,b;z)\begin{cases}(z\overline{\eta}-\alpha_{b})u(b)-\rho_{b}u(b-1),\quad&b\text{ even,}\\ (z\alpha_{b}-\eta)u(b)+z\rho_{b-1}u(b-1),&b\text{ odd}\end{cases}$
@@
Let $x,\omega\in\mathbb{T}^{d}$, $z\in\partial\mathbb{D}$ and $[a,b]\subset\mathbb{Z}$. If for any $m\in[a+2,b-2]$, there exists an interval $I_{m}=[a_{m},b_{m}]\subset[a+2,b-2]$ containing $m$ such that
$|G^{\beta,\eta}_{I_{m}}(a_{m},m;z)|\begin{cases}|z\overline{\beta}-\alpha_{a}|+\rho_{a},&a\text{ even,}\\ |z\alpha_{a}-\beta|+\rho_{a},&a\text{ odd,}\end{cases}+|G^{\beta,\eta}_{I_{m}}(m,b_{m};z)|\begin{cases}|z\overline{\eta}-\alpha_{b}|+\rho_{b},&b\text{ even}\\ |z\alpha_{b}-\eta|+\rho_{b-1},&b\text{ odd}\end{cases}<1,$
@@
Assume to the contrary that $z\in\sigma(\mathcal{E}^{\beta,\eta}_{[a,b]})$ and let $u$ be a corresponding eigenvector, i.e., [$\mathcal{E}^{\beta,\eta}_{[a,b]}u=zu$. Suppose that $|u(m)|=\max\limits_{n\in(a,b)}|u(n)|$. If $m\in[a+2,b-2]$, then there is a vector $\tilde{u}$ satisfying [$\mathcal{E}^{\beta,\eta}_{[a,b]}\tilde{u}=z\tilde{u}$ and $\tilde{u}|_{[a+2,b-2]}=u|_{[a+2,b-2]}$. Thus, from the hypothesis and the Poisson formula, one can get that
@@
Remark 3.24.
In the above covering lemma, the choice of the interval $[a+2,b-2]$ depends on the structure of the matrix [$\mathcal{E}^{\beta,\eta}_{[a,b]}$. Both this matrix and Poisson formula depend on the parity of $a$ and $b$. In addition, the purpose of the boundary condition $\beta$ and $\eta$ is to guarantee that the restriction of matrix $\mathcal{E}$ to $[a,b]$ is unitary. In fact, no matter how we choose $a$ and $b$ (the parity of $a$ and $b$), there are at most the first two rows and the last two rows relevant to the parameters $\beta$ and $\eta$. But under the condition of parameters $\beta$ and $\eta$, Poisson formula is no longer applicable. Thus, by considering the interval $[a+2,b-2]$ and applying the Poisson formula, we can prove the above covering lemma.
@@
Thus, according to Lemma 3.10, for $10\leq j\leq k\leq n-$ 1, one can obtain

### v3

@@ 3.2 Estimation of the Lyapunov Exponent
The Lyapunov exponent is a core indicator characterizing the spectral properties and localization behavior of operators. First, we need to clarify the deviation estimate between the finite scale approximation of the Lyapunov exponent $L_{n}(\omega,z)$ and its limit $L(\omega,z)$.
@@
It is straightforward to verify that to verify that $M(\omega,z;x)$ is conjugate to an $\mathbb{SL}(2,\mathbb{R})$ matrix $T(\omega,z;x)$ via
$Q=-\frac{1}{1+i}\left(\begin{array}[]{cc}1&-i\\ 1&i\end{array}\right)\in\mathbb{U}(2),$
@@
$T(\omega,z;x)=Q^{*}M(\omega,z;x)Q\in\mathbb{SL}(2,\mathbb{R}),$
@@
We can finally prove this lemma by summing over $2^{k}n$ . ∎
Based on the continuity result in the above lemma, we can further refine the stability of the logarithm of the transfer matrix norm under small perturbations.
@@
$M_{k,n}=\frac{1}{\rho(x_{k}+n\omega_{k})}\left(\begin{array}[]{cc}\sqrt{z_{k}}&-\frac{\overline{\alpha}(x_{k}+n\omega_{k})}{\sqrt{z_{k}}}\\ -\alpha(x_{k}+n\omega_{k})&1/\sqrt{z_{k}}\end{array}\right),\quad k=1,2.$
@@
It is not difficult to check that $M_{k,n}$ is conjugate to an $\mathbb{SL}(2,\mathbb{R})$ matrix $T_{k,n}$ through $Q$, that is,
@@
According to the polar decomposition of $\mathbb{SL}(2,\mathbb{C})$ matrices,
@@
[…]
@@
[…]
@@
[…]
@@
Now we study continuity from the transfer matrix level to the Lyapunov exponent level.
@@
Now we focus on the local continuity of the Lyapunov exponent with respect to frequency and spectral parameters.
@@
We further extend to the case of perturbations in the imaginary part of the phase with the help of the integral properties of subharmonic functions.
@@
[13, Lemma 4.1] Let $0<\rho<1$ and suppose $f$ is subharmonic on
@@
such that $\sup\limits_{\mathcal{A}_{\rho}}f\leq 1$ and $\int_{\mathbb{T}}f(x)dx\geq 0$. Then for any $y$, $y^{\prime}$ so that $\frac{\rho}{2}<y,y^{\prime}<\frac{\rho}{2}$ one has
$\Big|\int_{\mathbb{T}}f(x+iy)dx-\int_{\mathbb{T}}f(x+iy^{\prime})dx\Big|\leq C_{\rho}|y-y^{\prime}|.$
@@
$=C(z)\Big|\int_{\mathbb{T}}f(x_{1}+iy_{1};y_{2})dx_{1}-\int_{\mathbb{T}}f(x_{1};0)dx_{1}\Big|$
$=C(z)\Big|\int_{\mathbb{T}}(f(x_{1}+iy_{1};y_{2})-f(x_{1}+iy_{1};0))dx_{1}+\int_{\mathbb{T}}(f(x_{1}+iy_{1};0)-f(x_{1};0))dx_{1}\Big|$
$\leq C(z)\Big(\int_{\mathbb{T}}|f(x_{1}+iy_{1};y_{2})-f(x_{1}+iy_{1};0)|dx_{1}+\int_{\mathbb{T}}|f(x_{1}+iy_{1};0)-f(x_{1};0)|dx_{1}\Big)$
@@
The following lemma establishes the upper bound estimate of the logarithm of the transfer matrix norm.
@@
Lemma 3.10 establishes the global upper bound of the logarithm of the transfer matrix norm, and the following corollary extends this result to the case of small perturbations.
@@
Then we can further obtain the norm estimate of the difference between transfer matrices.
@@ 3.3 LDT Estimate for the Entries of the Monodromy
$\Theta_{n}=\left(\begin{array}[]{cc}\overline{\alpha}_{n}&\rho_{n}\\ \rho_{n}&-\alpha_{n}\end{array}\right).$
@@
$\mathcal{L}_{+}=\left(\begin{matrix}\Theta_{0}&~&~\\ ~&\Theta_{2}&~\\ ~&~&\ddots\end{matrix}\right),\quad\mathcal{M}_{+}=\left(\begin{matrix}\mathbf{1}&~&~\\ ~&\Theta_{1}&~\\ ~&~&\ddots\end{matrix}\right).$
@@
Without loss of generality, we restrict the extended CMV matrix to the interval [0,n1][0,n-1]. We denote the characteristic determinant of matrix $\mathcal{E}^{\beta,\eta}_{[0,n-1]}$ as $\varphi^{\beta,\eta}_{[0,n-1]}(\omega,z;x)=\det\big(z-\mathcal{E}^{\beta,\eta}_{[0,n-1]}\big)$. According to the results in [26, Theorem 2], the relation between this characteristic determinant and the $n$-step transfer matrix is
@@
To ensure the characteristic determinant does not vanish on a large set, we first establish a lower bound on its measure of small values using properties of real-analytic functions and matrix perturbations.
@@
Extending the previous result, we generalize the lower bound to smaller thresholds of the characteristic determinant, ensuring robustness across a range of small values that are relevant for LDT.
@@
Let $N_{0}$ be as in Lemma 3.15 and $0<\theta<1$. Fix some $z$ and let $\varphi_{[0,n-1]}^{\beta,\eta}=\det\big(z-\mathcal{E}^{\beta,\eta}_{[0,n-1]}\big)$. Then for any $N_{0}\leq n<N^{\frac{1-\theta}{2}}$,
@@
Now we introduce the properties of matrices in $\mathbb{SL}(2,\mathbb{C})$. For any $A\in\mathbb{SL}(2,\mathbb{C})$, there are unit vectors $\underline{u}^{+}_{A}$, $\underline{u}^{-}_{A}$, $\underline{v}^{+}_{A}$, $\underline{v}^{-}_{A}$ such that $A\underline{u}^{+}_{A}=\|A\|\underline{v}^{+}_{A}$, $A\underline{u}^{-}_{A}=\|A\|^{-1}\underline{v}^{-}_{A}$. Moreover, $\underline{u}^{+}_{A}\bot\underline{u}^{-}_{A}$ and $\underline{v}^{+}_{A}\bot\underline{v}^{-}_{A}$. In what follows, we denote that $a\wedge b:=\min\{a,b\}$ for positive numbers $a,b$ and $\varphi\wedge\psi$ stands for the exterior product of vectors $\varphi,\psi$.
@@
From the Theorem 3.2, we have
@@
From (3.19), the transfer matrix $M_{n}(\omega,z;x)\in\mathbb{S}\mathbb{U}(1,1)$, then we can verify that
$T_{n}(\omega,z;x)=Q^{*}M_{n}(\omega,z;x)Q=(\sqrt{z})^{-n}\Big(\prod_{j=0}^{n-1}\frac{1}{\rho_{j}}\Big)\left(\begin{matrix}t_{1}&t_{2}\\ t_{3}&t_{4}\end{matrix}\right)\in\mathbb{SL}(2,\mathbb{R}),$ (3.26)
@@
$\mathcal{P}_{n-1}=\left(\begin{array}[]{ccccccc}-\overline{\alpha}_{1}\rho_{0}&-\rho_{1}\rho_{0}&&&&\\ -\overline{\alpha}_{2}\rho_{1}&z+\overline{\alpha}_{2}\alpha_{1}&-\overline{\alpha}_{3}\rho_{2}&-\rho_{3}\rho_{2}&&\\ -\rho_{2}\rho_{1}&\rho_{2}\alpha_{1}&z+\overline{\alpha}_{3}\alpha_{2}&\rho_{3}\alpha_{2}&&\\ &&-\overline{\alpha}_{4}\rho_{3}&z+\overline{\alpha}_{4}\alpha_{3}&&\\ &&&&\ddots&\\ &&&&&z+\overline{\eta}\alpha_{n-2}\end{array}\right).$
@@
Next, we establish a lower bound on the average of the logarithm of the characteristic determinant, ensuring it is close to the finite scale Lyapunov exponent.
@@
Let $u=\frac{1}{n}\log|\varphi^{\beta,\eta}_{[0,n-1]}(x)|$, $\langle u\rangle:=\int_{\mathbb{T}^{2}}u(x)dx$. From (3.28), (3.29) and (3.19), one can get
@@
Let $g(z)=\int_{\mathbb{T}}u(z,y)dy$. Then $\sup\limits_{\mathbb{T}}g<\sup\limits_{\mathbb{T}^{2}}u<L_{n}+Cn^{-\tau}$ and $\langle g\rangle=\langle u\rangle>L_{n}-n^{-\kappa}$. This implies that
@@
$\|(\mathcal{E}^{\beta,\eta}_{[0,n-1]}(\omega,z;x)-z)^{-1}\|\geq C\exp(H+n^{\nu}/d)$
@@
Due to the structure of the matrix $\mathcal{E}^{\beta,\eta}_{[0,n-1]}$, there exists some constant $C$ such that
$\|\mathcal{E}^{\beta,\eta}_{[0,n-1]}(\omega,z;x)-\mathcal{E}^{\beta,\eta}_{[0,n-1]}(\omega,z;x^{\prime})\|\leq C|x-x^{\prime}|.$
@@
$\|(\mathcal{E}^{\beta,\eta}_{[0,n-1]}(\omega,z;x^{\prime})-E)^{-1}\|\|\mathcal{E}^{\beta,\eta}_{[0,n-1]}(\omega,z;x)-\mathcal{E}^{\beta,\eta}_{[0,n-1]}(\omega,z;x^{\prime})\|<1,$
@@
$\|(\mathcal{E}^{\beta,\eta}_{[0,n-1]}(\omega,z;x^{\prime})-E)^{-1}\|\|\mathcal{E}^{\beta,\eta}_{[0,n-1]}(\omega,z;x)-\mathcal{E}^{\beta,\eta}_{[0,n-1]}(\omega,z;x^{\prime})\|$
$=\|(\mathcal{E}^{\beta,\eta}_{[0,n-1]}(\omega,z;x^{\prime})-E)^{-1}\|\|(\mathcal{E}^{\beta,\eta}_{[0,n-1]}(\omega,z;x)-E)-(\mathcal{E}^{\beta,\eta}_{[0,n-1]}(\omega,z;x^{\prime})-E)\|$
@@
Thus, $E-[E-(\mathcal{E}^{\beta,\eta}_{[0,n-1]}(\omega,z;x^{\prime})-E)^{-1}(\mathcal{E}^{\beta,\eta}_{[0,n-1]}(\omega,z;x)-E)]$ is invertible and this implies that $\mathcal{E}^{\beta,\eta}_{[0,n-1]}(\omega,z;x)-E$ is invertible. Then the second statement follows. ∎
@@ 3.4 Poisson Formula for CMV Matrix
$u(m)=$ $G^{\beta,\eta}_{[a,b]}(a,m;z)\begin{cases}(z\overline{\beta}-\alpha_{a})u(a)-\rho_{a}u(a+1),\quad&a\text{ even}\\ (z\alpha_{a}-\beta)u(a)+z\rho_{a}u(a+1),&a\text{ odd}\end{cases}$
$+G^{\beta,\eta}_{[a,b]}(m,b;z)\begin{cases}(z\overline{\eta}-\alpha_{b})u(b)-\rho_{b}u(b-1),\quad&b\text{ even}\\ (z\alpha_{b}-\eta)u(b)+z\rho_{b-1}u(b-1),&b\text{ odd}\end{cases}$
@@
Let $x,\omega\in\mathbb{T}^{d}$, $z\in\partial\mathbb{D}$ and $a,b]\subset\mathbb{Z}$. If for any $m\in[a+1,b-1]$, there exists an interval $I_{m}=[a_{m},b_{m}]\subset[a+1,b-1]$ containing $m$ such that
$|G^{\beta,\eta}_{I_{m}}(a_{m},m;z)|\begin{cases}|z\overline{\beta}-\alpha_{a}|+\rho_{a},&a\text{ even}\\ |z\alpha_{a}-\beta|+\rho_{a},&a\text{ odd}\end{cases}+|G^{\beta,\eta}_{I_{m}}(m,b_{m};z)|\begin{cases}|z\overline{\eta}-\alpha_{b}|+\rho_{b},&b\text{ even}\\ |z\alpha_{b}-\eta|+\rho_{b-1},&b\text{ odd}\end{cases}<1,$
@@
Assume to the contrary that $z\in\sigma(\mathcal{E}^{\beta,\eta}_{[a,b]})$ and let $u$ be a corresponding eigenvector, i.e., $\mathcal{E}^{\beta,\eta}_{[a,b]}u=zu$. Suppose that $|u(m)|=\max\limits_{n\in(a,b)}|u(n)|$. If $m\in[a+1,b-1]$, then there is a vector $\tilde{u}$ satisfying $\mathcal{E}^{\beta,\eta}_{[a,b]}\tilde{u}=z\tilde{u}$ and $\tilde{u}|_{[a+1,b-1]}=u|_{[a+1,b-1]}$. Thus, from the hypothesis and the Poisson formula, one can get that
@@
Next, we use the LDT for the characteristic determinant to bound Green’s functions and their operator norms.
@@
Thus, according to Lemma 3.10, for $0\leq j\leq k\leq n-1$, one can obtain